OpenStudy (anonymous):

Solve t = p+3r/6 for r.

OpenStudy (turingtest):

answer or process?

OpenStudy (turingtest):

r=2(t-p)

OpenStudy (turingtest):

unless you forgot parentheses somewhere...

jhonyy9 (jhonyy9):

t=p+3r/6 6t=6p+3r 6t-6p=3r r=6(t-p)/3 r=2(t-p)